For two events, E and F, you are given:
P(E) = 0.30
P(F) = 0.45
E and F are mutually exclusive
What is the value of P(E or F)? Enter your answer as a decimal rounded to two decimal places.

Answer by ikleyn(52813)

.

Since the events E and F are mutually exclusive, they are disjoint (have empty intersection).


Therefore, P(E or F) = P(E) + P(F) = 0.30 + 0.45 = 0.75.    ANSWER

Solved.